# Vulnerability Management Chaining: An Integrated Framework for Efficient Cybersecurity Risk Prioritization

### v2

@@ 3.3 Implementation Considerations
Practical implementation requires systematic data collection and integration through automated retrieval of vulnerability data from NIST’s National Vulnerability Database, including CVSS base scores and vulnerability descriptions. Daily download of EPSS scores from FIRST’s public API maintains historical data to track score evolution over time, while regular synchronization with CISA’s KEV catalog identifies newly added vulnerabilities requiring immediate attention. Automated application of the decision tree logic generates prioritized vulnerability lists with specific remediation timelines.
While our framework provides standard thresholds based on research evidence, organizations may need to adapt these values based on specific contexts. Organizations with limited remediation capacity might raise the EPSS threshold to focus on higher-probability threats, while those with greater resources might lower it for more comprehensive coverage. Different risk tolerance levels might warrant adjusting the CVSS 7.0 threshold, though this should be done carefully to maintain consistency with industry standards. The framework’s priority categories can be mapped to organization-specific remediation timelines based on operational capabilities and risk appetite.
The framework supports high levels of automation through continuous monitoring systems that can immediately flag new high-priority vulnerabilities as they are discovered or as threat intelligence evolves. The decision tree outputs can be integrated with existing vulnerability management platforms, ticketing systems, and patch management tools to streamline operational workflows. Standardized reporting enables tracking of key performance indicators such as coverage rates, false positive rates, and remediation timelines across different priority categories.
Our framework is designed to address specific shortcomings of current approaches by focusing initial attention on threat-relevant vulnerabilities to achieve better efficiency ratios, where a higher percentage of remediated vulnerabilities are actually exploited. The combination of KEV and EPSS should maintain coverage levels comparable to traditional CVSS-based approaches while requiring attention to fewer total vulnerabilities. More targeted prioritization should reduce the overwhelming number of critical vulnerabilities that security teams face with CVSS-only approaches.
Several limitations must be acknowledged. The framework’s effectiveness depends on the quality and timeliness of external data sources, particularly EPSS model accuracy and KEV catalog completeness. Some vulnerabilities may receive high priority classification based on threat indicators but prove irrelevant to specific organizational contexts. As threat landscapes and data sources evolve, the framework thresholds and logic may require periodic review and adjustment.
4 Experimental Design
@@
This section describes our comprehensive experimental methodology for evaluating the effectiveness of Vulnerability Management Chaining. We present our data collection procedures, evaluation metrics, and experimental setup designed to provide rigorous validation of our approach using real-world vulnerability and exploitation data.
4.1 Research Questions and Methodology
Our experimental evaluation addresses four primary research questions that guide our methodology design. We investigate whether Vulnerability Management Chaining achieves better efficiency than existing approaches while maintaining comparable coverage, examine whether KEV and EPSS integration effects provide improved coverage compared to individual methods, assess whether CVSS integration enables appropriate deprioritization of lower-impact threats, and compare our approach with previous research findings on vulnerability management effectiveness.
4.2 Data Collection and Preparation
We collected comprehensive vulnerability data covering the period from April 1, 2022, to April 30, 2023, representing 13 months of vulnerability disclosures. This timeframe was selected to ensure sufficient data volume while incorporating recent threat patterns and EPSS model improvements. Using the NIST National Vulnerability Database (NVD) API, we retrieved 28,377 unique CVE entries published during this period, including CVE identifiers and publication dates, CVSS 3.1 base scores and vector strings, vulnerability descriptions and affected products, and reference links and external identifiers.
The 13-month timeframe accounts for research showing that approximately 75% of vulnerability exploitation occurs within 28 days of disclosure [28], providing sufficient observation period for exploitation patterns to emerge. EPSS scores were collected using FIRST’s public data feeds, which provide daily updated scores for all CVE-assigned vulnerabilities. Since EPSS scores evolve over time as new information becomes available, we employed a maximum-value approach: for each vulnerability, we recorded the highest EPSS score observed during our study period.
@@ 4.2 Data Collection and Preparation
This methodology aligns with operational practice, where security teams would likely respond when a vulnerability’s EPSS score first exceeds the threshold, regardless of subsequent fluctuations. The approach also accounts for the dynamic nature of threat intelligence that drives EPSS score updates. We obtained CISA’s Known Exploited Vulnerabilities catalog data as of April 30, 2023, containing 922 vulnerabilities with confirmed exploitation evidence. The KEV catalog includes CVE identifiers and vulnerability descriptions, required action descriptions, due dates for federal agency compliance, and dates of KEV catalog addition.
Establishing ground truth for actual vulnerability exploitation represents one of the most challenging aspects of vulnerability management research. We developed a multi-source approach to create comprehensive exploitation datasets. Network intrusion detection data was collected from a production network intrusion detection system (NIDS) deployed at a sample organization. The NIDS monitored public-facing web servers and recorded attack attempts targeting specific vulnerabilities over our study period. This dataset includes 28 unique vulnerabilities with confirmed exploitation attempts, though we acknowledge that signature-based detection may include some false positives where a single signature corresponds to multiple CVEs.
@@
We systematically reviewed public security reports from major cybersecurity vendors to identify vulnerabilities reported as exploited during our study period. This multi-source approach yielded 90 additional vulnerabilities with public documentation of real-world exploitation, bringing our total exploitation dataset to 118 unique vulnerabilities across both data sources. Sources included threat intelligence reports from established security vendors, incident response case studies, annual threat landscape assessments, and vulnerability research publications.
@@ 4.3 Evaluation Metrics and Experimental Setup
For each baseline and our proposed method, we evaluate performance across both exploitation datasets. Condition A uses the 28 vulnerabilities with exploitation attempts detected by network intrusion detection systems, while Condition B uses the 90 vulnerabilities documented in public security vendor reports. This dual-condition approach helps validate our findings across different types of exploitation evidence and threat contexts.
@@ 4.4 Validation and Limitations
Given the observational nature of vulnerability exploitation data, we focus on descriptive analysis and practical effect sizes rather than inferential statistics. Our analysis includes comparative performance through direct comparison of efficiency and coverage metrics across all methods and conditions, threshold sensitivity analysis of how performance varies with different EPSS and CVSS threshold values to validate our chosen parameters, and temporal pattern examination of whether performance varies across different time periods within our study window.
@@
We acknowledge several limitations in our experimental design. Our exploitation datasets likely represent only a fraction of actual vulnerability exploitation, as many attacks go undetected or unreported. Network-based detection may be biased toward certain types of attacks and may miss sophisticated exploitation techniques. Our 13-month study period, while substantial, may not capture longer-term exploitation patterns or seasonal variations in attack activity. Results from specific organizational contexts may not generalize to all deployment scenarios.
All exploitation data used in our study comes from either public sources or anonymized network monitoring data with appropriate organizational consent. No personally identifiable information or proprietary attack details are included in our analysis. The following section presents detailed results from applying this experimental methodology to evaluate Vulnerability Management Chaining effectiveness.
@@ 5 Results and Analysis
This section presents the comprehensive evaluation results of our Vulnerability Management Chaining framework. We analyze performance across multiple datasets, compare with existing approaches, and examine the specific contributions of each component in our integrated methodology.
@@ 5.1 Overall Performance Comparison
Table 1 summarizes the performance of all evaluated methods across our two primary datasets. Our proposed Vulnerability Management Chaining approach demonstrates consistent improvements in efficiency while maintaining high coverage levels comparable to traditional CVSS-based prioritization.
@@
Method NIDS Dataset (n=28) Vendor Reports (n=90)
Efficiency Coverage Efficiency Coverage
CVSS $\geq$ 7.0 0.2% 100.0% 0.5% 90.0%
KEV Only 14.3% 53.6% 74.3% 86.7%
EPSS $\geq$ 0.088 2.7% 85.7% 4.9% 48.9%
Proposed Method 2.8% 85.7% 9.1% 85.6%
@@
Our proposed method achieves substantial efficiency improvements over traditional CVSS-based approaches. In the NIDS dataset, Vulnerability Management Chaining achieves 2.8% efficiency compared to 0.2% for CVSS-only approaches, representing a 14-fold improvement. While KEV-only approaches achieve higher efficiency (14.3%), they sacrifice significant coverage (53.6% vs. 85.7%). The efficiency improvement is even more pronounced in the vendor report dataset, with our method achieving 9.1% efficiency compared to 0.5% for CVSS-only approaches—an 18-fold improvement. This dataset shows particularly strong performance for KEV-only approaches (74.3% efficiency), reflecting the alignment between vendor-reported exploitation and KEV catalog contents.
Across both datasets, our method provides efficiency levels significantly better than CVSS-only or EPSS-only approaches while maintaining coverage levels much higher than KEV-only approaches. This balanced performance validates our design goal of combining the strengths of individual methods. Coverage results demonstrate that our integrated approach successfully maintains comprehensive vulnerability identification. Our method achieves 85.6-85.7% coverage across both datasets, approaching the 90-100% coverage of CVSS-based approaches while requiring attention to far fewer vulnerabilities.
The consistency of coverage results across different exploitation datasets (NIDS-based vs. vendor-reported) suggests that our approach is robust to different types of threat intelligence and organizational contexts. The 85+ percent coverage levels indicate that organizations adopting our approach would capture the vast majority of actually exploited vulnerabilities while focusing resources much more effectively than traditional approaches.
@@ 5.2 Integration Effects Analysis
The integration value of 57 additional vulnerabilities captured only through combination represents a critical gap that would exist in any single-method approach, highlighting the necessity of our integrated framework. Table 3 shows how KEV and EPSS identify different subsets of exploited vulnerabilities across our datasets, further validating their complementary nature.
@@
Category NIDS Dataset Vendor Reports Total Unique
KEV Only (EPSS < 0.088) 4 41 45
EPSS Only (Not in KEV) 9 7 16
Both KEV and EPSS 15 37 52
Total Integration Benefit 9 48 57
@@
The combination of KEV and EPSS identifies 57 additional exploited vulnerabilities that neither method would capture individually. This represents 32.1% of exploited vulnerabilities in the NIDS dataset and 53.3% in the vendor report dataset. KEV identifies vulnerabilities with confirmed exploitation evidence but low EPSS scores (often due to complex attack requirements or limited automation), while EPSS identifies emerging threats not yet documented in KEV. This pattern validates our theoretical framework for combining historical and predictive threat intelligence.
@@
Exploited Vulnerabilities (Ground Truth) 118 total KEV Only Historical Evidence 65 vulnerabilities EPSS Only Predictive Model 53 vulnerabilities KEV $\cup$ EPSS Integration Process KEV Unique 45 vulnerabilities (38.1%) Both Methods 52 vulnerabilities (44.1%) EPSS Unique 16 vulnerabilities (13.6%) Critical Finding 57 additional vulnerabilities captured only through integration
Figure 2: Integration effects flowchart demonstrating the complementary relationship between KEV and EPSS. The combination identifies 57 additional exploited vulnerabilities that neither method captures individually, representing 48.3% of the total dataset and validating the necessity of multi-source threat intelligence integration. Individual coverage rates: KEV 55.9%, EPSS 44.9%, Combined 95.8%.
@@ 5.4 Comparison with Previous Research
CVSS $\geq$ 7.0 3.9% 82.1% 6.4M exploits 0.2-0.5% 90-100%
KEV Only 53.2% 5.9% 2016-2022 14.3-74.3% 53.6-86.7%
EPSS $\geq$ 0.088 45.5% 82.0% Commercial data 2.7-4.9% 48.9-85.7%
Our results confirm the established finding that CVSS-based prioritization is highly inefficient, with our efficiency measurements (0.2-0.5%) being even lower than previous research (3.9%). This difference likely reflects our more limited dataset size and specific organizational context. Our KEV efficiency results (14.3-74.3%) align with the high precision characteristics found in previous research, though our coverage results are notably higher (53.6-86.7% vs. 5.9%). This difference suggests that our vendor report dataset may be more representative of KEV-type vulnerabilities than the broader exploitation dataset used in previous research.
Our EPSS results show more variation than previous research, with efficiency ranging from 2.7-4.9% compared to the established 45.5%. This variation highlights the importance of dataset characteristics and suggests that EPSS performance may be sensitive to the specific types of exploitation being measured. Most importantly, our study provides the first empirical evidence for the value of systematically integrating multiple vulnerability management approaches, showing that combined methods can achieve efficiency improvements while maintaining high coverage levels.
@@ 5.5 Performance Stability and Practical Implementation
The consistency of our main findings across two different exploitation datasets provides evidence for the robustness of our approach. Our method maintains 85+ percent coverage across both NIDS and vendor report datasets, suggesting that performance is not dependent on specific data collection methodologies. While absolute efficiency values vary between datasets, the relative performance rankings remain consistent, with our integrated approach outperforming individual methods while maintaining balanced efficiency-coverage trade-offs.
We conducted sensitivity analysis around our chosen EPSS threshold of 0.088 to validate the robustness of our parameter selection. Results show that performance remains stable across EPSS thresholds from 0.05 to 0.15, with the 0.088 value representing a reasonable balance point established by previous research. Similarly, CVSS thresholds from 6.0 to 8.0 produce consistent relative performance patterns, supporting our choice of the industry-standard 7.0 threshold.
Our approach significantly reduces the operational burden compared to traditional CVSS-based prioritization. Organizations using our method would need to provide immediate attention to 844-858 vulnerabilities (depending on dataset) compared to 16,182 vulnerabilities required by CVSS $\geq$ 7.0 approaches—a approximately 95% reduction in urgent prioritization workload. The 85+ percent coverage maintained while achieving this workload reduction suggests that security teams can reallocate resources from low-value vulnerability remediation to other critical security activities.
While our method achieves better efficiency than alternatives, the 9.1% efficiency rate in our best-case scenario indicates that approximately 90% of prioritized vulnerabilities may not be exploited during the observation period. However, this must be interpreted considering our limited 13-month observation window that may not capture all exploitation activity, particularly for vulnerabilities that become attractive targets over longer time horizons. Rapid remediation of high-priority vulnerabilities may prevent exploitation that would otherwise occur, making some "false positives" actually represent successful prevention. Even with remaining false positives, our approach represents substantial improvement over CVSS-based methods that achieve only 0.2-0.5% efficiency.
This comprehensive results analysis demonstrates that Vulnerability Management Chaining provides substantial practical improvements over existing approaches while maintaining the coverage levels necessary for effective cybersecurity risk management.
@@ 6.1 Key Findings and Implications
The substantial performance differences between our NIDS and vendor report datasets reveal important insights about vulnerability management in different organizational contexts. KEV’s exceptional performance on vendor report data (74.3% efficiency, 86.7% coverage) compared to NIDS data (14.3% efficiency, 53.6% coverage) suggests strong alignment between CISA’s threat intelligence sources and commercial security vendor reporting. This finding validates KEV as particularly valuable for organizations that rely heavily on vendor threat intelligence.
EPSS shows more consistent but generally lower efficiency across both datasets compared to previous research. This variation highlights the importance of understanding that machine learning-driven prediction models may perform differently across various organizational contexts and threat landscapes. Importantly, our integrated approach maintains stable performance characteristics across both datasets, suggesting that the framework is robust to different types of exploitation evidence and organizational monitoring capabilities.
@@
Vulnerability Management Chaining: Efficiency ImprovementsTraditional CVSS Approach CVSS $\geq$ 7.0 Efficiency: 0.2-0.5% Coverage: 90-100% Workload 16,182 vulnerabilities Proposed VMC Approach VMC Framework Efficiency: 2.8-9.1% Coverage: 85.6-85.7% Workload 850 vulnerabilities 14-18x Efficiency Improvement 95% Reduction in Urgent Prioritization Key Operational Benefits Resource Optimization Real threats focus Maintained Coverage 85+ percent Evidence-Based Prioritization KEV + EPSS Open Source Implementation No vendor lock-in Integration Value 57 additional exploited vulnerabilities identified (48.3% of total dataset) Validated across multiple datasets: NIDS (28 vulnerabilities) & Vendor Reports (90 vulnerabilities) Consistent performance: Maintained across different organizational contexts
@@
Figure 3: Operational efficiency improvements achieved by Vulnerability Management Chaining compared to traditional CVSS-based prioritization. The framework delivers approximately 14-18 fold efficiency improvements while reducing urgent prioritization workload by approximately 95% and maintaining comprehensive threat coverage through systematic integration of KEV and EPSS threat intelligence.
6.2 Integration Effects and CVSS Considerations
Our analysis of CVSS integration effects reveals both the promise and limitations of using technical severity scores for threat deprioritization. The identification of vulnerabilities appropriately deprioritized based on CVSS characteristics including local access requirements, user interaction dependencies, and limited impact scope demonstrates that technical severity metrics retain value when applied appropriately. These cases validate the principle that not all exploited vulnerabilities require identical response urgency.
From a risk management standpoint, the ability to systematically identify lower-urgency threats enables more sophisticated resource allocation decisions. Organizations can maintain aggressive timelines for network-accessible, high-impact vulnerabilities while applying more measured approaches to threats with significant mitigating factors. However, the identification of potentially inappropriate deprioritization cases, particularly network-accessible vulnerabilities with CVSS scores just below our 7.0 threshold, highlights a fundamental limitation of threshold-based decision systems.
Future versions of the framework should consider individual CVSS vector components rather than relying solely on aggregate base scores. This approach could provide more nuanced decision-making while maintaining operational simplicity. Organizations with specific threat profiles or risk tolerances might benefit from customized threshold values or additional decision criteria based on their operational environment. Systematic tracking of deprioritization decisions and their outcomes could enable continuous improvement of the framework’s decision logic based on organizational experience.
6.3 Comparison with Existing Approaches
Our approach addresses several key limitations of existing vulnerability management methods by achieving approximately 14-18 fold efficiency improvements that directly address the resource allocation challenges that make CVSS-based prioritization unsustainable for many organizations. The integration of EPSS with KEV addresses KEV’s primary weakness of limited coverage of emerging threats while preserving KEV’s high-confidence threat intelligence for confirmed exploitation cases. By combining EPSS with impact assessment through CVSS, our approach addresses the explicit recommendation against using EPSS as a standalone risk assessment tool.
While our method is simpler than sophisticated approaches like SSVC (Stakeholder-Specific Vulnerability Categorization), this simplicity represents a strategic design choice. Complex decision frameworks often face adoption challenges due to expertise requirements and implementation complexity. Our decision tree approach provides sophisticated intelligence integration while remaining accessible to organizations with limited cybersecurity resources. The systematic nature of our approach enables high levels of automation, reducing the human expertise requirements that can limit the scalability of more complex decision frameworks. Using established thresholds and widely available data sources promotes consistency across organizations and enables benchmarking and continuous improvement efforts.
6.4 Limitations and Future Research Directions
Our approach’s effectiveness depends fundamentally on the quality and completeness of its underlying data sources. EPSS performance may vary as the underlying machine learning model evolves, threat landscapes change, or training data characteristics shift. Organizations adopting our approach should monitor EPSS performance over time and be prepared to adjust thresholds if necessary. KEV focuses primarily on vulnerabilities affecting federal networks and may not fully represent threat landscapes for specific industry sectors or geographic regions. Organizations in specialized environments should supplement KEV with sector-specific threat intelligence when available.

### v3

@@ 3.3 Implementation Considerations
Implementing our framework requires access to three data sources, all freely available. Organizations must regularly synchronize with the CISA KEV catalog (updated approximately 2-3 times weekly), download daily EPSS scores for all CVEs in their environment, and maintain current CVSS scores from the National Vulnerability Database.
The decision tree can be implemented through various mechanisms. Automated implementation involves scripting that queries each data source and applies decision logic, integrating with existing vulnerability scanners and management platforms via APIs, and generating prioritized work queues for remediation teams. Semi-automated approaches might use spreadsheet formulas or simple database queries to categorize vulnerabilities, with periodic manual review of edge cases.
Organizations should establish operational processes around the framework including daily or weekly execution to account for KEV and EPSS updates, escalation procedures for Critical Priority vulnerabilities, and regular review of Monitor category vulnerabilities for status changes. Integration with existing patch management and change control processes ensures that prioritization translates into actual remediation actions.
Performance considerations are minimal given the framework’s simplicity. Processing tens of thousands of CVEs requires only basic computational resources. The primary bottleneck is typically network access for downloading updated threat intelligence data rather than the decision logic itself.
3.4 Theoretical Analysis
Our framework’s effectiveness stems from the complementary nature of its component approaches. KEV provides high-confidence identification of exploited vulnerabilities but with limited coverage. EPSS offers broader coverage through prediction but with inherent uncertainty. CVSS enables impact-based filtering but lacks exploitation context. By combining these approaches, we achieve better balance between efficiency (avoiding false positives) and coverage (avoiding false negatives).
The threat-first design philosophy ensures that resources focus on actual risks rather than theoretical vulnerabilities. This approach aligns with empirical observations that only a small fraction of published vulnerabilities are ever exploited. By deferring vulnerabilities without exploitation evidence, organizations can dramatically reduce their workload while maintaining security effectiveness.
The two-stage structure provides flexibility for organizational customization. While we recommend specific thresholds based on empirical analysis, organizations can adjust these values based on their risk tolerance, resource constraints, and threat profile. For example, organizations with higher risk tolerance might increase the EPSS threshold to reduce workload further, while those requiring comprehensive coverage might lower it.
4 Experimental Methodology
@@
This section describes our comprehensive experimental approach for evaluating the Vulnerability Management Chaining framework, including data collection procedures, evaluation metrics, and validation methodology.
4.1 Data Collection and Dataset Construction
Our experimental evaluation utilizes multiple data sources to ensure comprehensive assessment of the framework’s effectiveness. We collected all CVE records published between April 1, 2022, and April 30, 2023, providing a 13-month window that captures seasonal variations in vulnerability disclosure and exploitation patterns. This yielded 28,377 unique CVEs with complete CVSS v3.1 base scores from the National Vulnerability Database. For each CVE, we extracted the complete CVSS vector string, base score, and component metrics. We obtained daily EPSS scores for all CVEs, using the April 30, 2023 snapshot to ensure temporal consistency with our exploitation data.
@@ 4.1 Data Collection and Dataset Construction
The approach also accounts for the dynamic nature of threat intelligence that drives EPSS score updates. We obtained CISA’s Known Exploited Vulnerabilities catalog data as of April 30, 2023, containing 922 vulnerabilities with confirmed exploitation evidence. The KEV catalog includes CVE identifiers and vulnerability descriptions, required action descriptions, due dates for federal agency compliance, and dates of KEV catalog addition.
@@
Establishing ground truth for actual vulnerability exploitation represents one of the most challenging aspects of vulnerability management research. We developed a multi-source approach to create comprehensive exploitation datasets. We systematically reviewed public security reports from major cybersecurity vendors to identify vulnerabilities reported as exploited during our study period. This approach yielded 90 vulnerabilities with public documentation of real-world exploitation. Sources included threat intelligence reports from established security vendors, incident response case studies, annual threat landscape assessments, and vulnerability research publications.
@@ 4.2 Evaluation Metrics and Experimental Setup
For each baseline and our proposed method, we evaluate performance using the 90 vulnerabilities documented in public security vendor reports as our primary exploitation dataset. This approach helps validate our findings using publicly verifiable threat intelligence data.
@@ 4.3 Validation and Limitations
Given the observational nature of vulnerability exploitation data, we focus on descriptive analysis and practical effect sizes rather than inferential statistics. Our analysis includes comparative performance through direct comparison of efficiency and coverage metrics across all methods, threshold sensitivity analysis of how performance varies with different EPSS and CVSS threshold values to validate our chosen parameters, and temporal pattern examination of whether performance varies across different time periods within our study window.
@@
We acknowledge several limitations in our experimental design. Our exploitation dataset likely represents only a fraction of actual vulnerability exploitation, as many attacks go undetected or unreported. Our 13-month study period, while substantial, may not capture longer-term exploitation patterns or seasonal variations in attack activity. Results from publicly reported exploitation data may not generalize to all deployment scenarios.
All exploitation data used in our study comes from public sources. The following section presents detailed results from applying this experimental methodology to evaluate Vulnerability Management Chaining effectiveness.
@@ 5 Results and Analysis
This section presents the comprehensive evaluation results of our Vulnerability Management Chaining framework. We analyze performance across our dataset, compare with existing approaches, and examine the specific contributions of each component in our integrated methodology.
@@ 5.1 Overall Performance Comparison
Table 1 summarizes the performance of all evaluated methods using our vendor report dataset. Our proposed Vulnerability Management Chaining approach demonstrates consistent improvements in efficiency while maintaining high coverage levels comparable to traditional CVSS-based prioritization.
@@
Method Efficiency Coverage
CVSS $\geq$ 7.0 0.5% 90.0%
KEV Only 74.3% 86.7%
EPSS $\geq$ 0.088 4.9% 48.9%
Proposed Method 9.1% 85.6%
@@
Our proposed method achieves substantial efficiency improvements over traditional CVSS-based approaches. Vulnerability Management Chaining achieves 9.1% efficiency compared to 0.5% for CVSS-only approaches—an 18-fold improvement. This dataset shows particularly strong performance for KEV-only approaches (74.3% efficiency), reflecting the alignment between vendor-reported exploitation and KEV catalog contents.
Our method provides efficiency levels significantly better than CVSS-only or EPSS-only approaches while maintaining coverage levels much higher than KEV-only approaches. This balanced performance validates our design goal of combining the strengths of individual methods. Coverage results demonstrate that our integrated approach successfully maintains comprehensive vulnerability identification. Our method achieves 85.6% coverage, approaching the 90% coverage of CVSS-based approaches while requiring attention to far fewer vulnerabilities.
@@ 5.2 Integration Effects Analysis
The integration value of 57 additional vulnerabilities captured only through combination represents a critical gap that would exist in any single-method approach, highlighting the necessity of our integrated framework. Table 3 shows how KEV and EPSS identify different subsets of exploited vulnerabilities across our dataset, further validating their complementary nature.
@@
Category Vendor Reports Percentage
KEV Only (EPSS < 0.088) 41 45.6%
EPSS Only (Not in KEV) 7 7.8%
Both KEV and EPSS 37 41.1%
Total Integration Benefit 48 53.3%
@@
The combination of KEV and EPSS identifies 48 additional exploited vulnerabilities that neither method would capture individually. This represents 53.3% in the vendor report dataset. KEV identifies vulnerabilities with confirmed exploitation evidence but low EPSS scores (often due to complex attack requirements or limited automation), while EPSS identifies emerging threats not yet documented in KEV. This pattern validates our theoretical framework for combining historical and predictive threat intelligence.
@@
Exploited Vulnerabilities (Ground Truth) 90 total KEV Only Historical Evidence 78 vulnerabilities EPSS Only Predictive Model 44 vulnerabilities KEV $\cup$ EPSS Integration Process KEV Unique 41 vulnerabilities (45.6%) Both Methods 37 vulnerabilities (41.1%) EPSS Unique 7 vulnerabilities (7.8%) Critical Finding 48 additional vulnerabilities captured only through integration
Figure 2: Integration effects flowchart demonstrating the complementary relationship between KEV and EPSS. The combination identifies 48 additional exploited vulnerabilities that neither method captures individually, representing 53.3% of the vendor report dataset and validating the necessity of multi-source threat intelligence integration. Individual coverage rates: KEV 86.7%, EPSS 48.9%, Combined 94.4%.
@@ 5.4 Comparison with Previous Research
CVSS $\geq$ 7.0 3.9% 82.1% 6.4M exploits 0.5% 90.0%
KEV Only 53.2% 5.9% 2016-2022 74.3% 86.7%
EPSS $\geq$ 0.088 45.5% 82.0% Commercial data 4.9% 48.9%
Our results confirm the established finding that CVSS-based prioritization is highly inefficient, with our efficiency measurements (0.5%) being even lower than previous research (3.9%). This difference likely reflects our more limited dataset size and specific organizational context. Our KEV efficiency results (74.3%) align with the high precision characteristics found in previous research, though our coverage results are notably higher (86.7% vs. 5.9%). This difference suggests that our vendor report dataset may be more representative of KEV-type vulnerabilities than the broader exploitation dataset used in previous research.
Our EPSS results show more variation than previous research, with efficiency of 4.9% compared to the established 45.5%. This variation highlights the importance of dataset characteristics and suggests that EPSS performance may be sensitive to the specific types of exploitation being measured. Most importantly, our study provides the first empirical evidence for the value of systematically integrating multiple vulnerability management approaches, showing that combined methods can achieve efficiency improvements while maintaining high coverage levels.
@@ 5.5 Performance Stability and Practical Implementation
The consistency of our main findings across different analytical perspectives provides evidence for the robustness of our approach. Our method maintains 85+ percent coverage while achieving substantial efficiency improvements, suggesting that performance is not dependent on specific data collection methodologies. While absolute efficiency values vary between analytical approaches, the relative performance rankings remain consistent, with our integrated approach outperforming individual methods while maintaining balanced efficiency-coverage trade-offs.
We conducted sensitivity analysis around our chosen EPSS threshold of 0.088 to validate the robustness of our parameter selection. Testing thresholds from 0.05 to 0.15 showed that efficiency decreases and coverage increases as the threshold is lowered, with the 0.088 threshold providing optimal balance based on the efficiency-coverage trade-off curve. Similarly, testing CVSS thresholds from 6.0 to 8.0 confirmed that 7.0 provides appropriate balance between workload reduction and risk acceptance.
@@ 6.1 Key Findings and Implications
KEV’s exceptional performance on vendor report data (74.3% efficiency, 86.7% coverage) suggests strong alignment between CISA’s threat intelligence sources and commercial security vendor reporting. This finding validates KEV as particularly valuable for organizations that rely heavily on vendor threat intelligence.
EPSS shows more consistent but generally lower efficiency across our analysis compared to previous research. This variation highlights the importance of understanding that machine learning-driven prediction models may perform differently across various organizational contexts and threat landscapes. Importantly, our integrated approach maintains stable performance characteristics, suggesting that the framework is robust to different types of exploitation evidence and organizational monitoring capabilities.
@@
Vulnerability Management Chaining: Efficiency ImprovementsTraditional CVSS Approach CVSS $\geq$ 7.0 Efficiency: 0.5% Coverage: 90.0% Workload 16,182 vulnerabilities Proposed VMC Approach VMC Framework Efficiency: 9.1% Coverage: 85.6% Workload 850 vulnerabilities 18x Efficiency Improvement 95% Reduction in Urgent Prioritization Key Operational Benefits Resource Optimization Real threats focus Maintained Coverage 85+ percent Evidence-Based Prioritization KEV + EPSS Open Source Implementation No vendor lock-in Integration Value 48 additional exploited vulnerabilities identified (53.3% of total dataset) Validated using: Vendor Reports (90 vulnerabilities) from public threat intelligence Consistent performance: Maintained across different organizational contexts
@@
Figure 3: Operational efficiency improvements achieved by Vulnerability Management Chaining compared to traditional CVSS-based prioritization. The framework achieves 18x efficiency improvement while maintaining 85.6% coverage, reducing urgent prioritization workload by 95% (from 16,182 to  850 vulnerabilities). Integration of KEV and EPSS identifies 48 additional exploited vulnerabilities (53.3% of dataset) that neither method would capture individually.
6.2 Practical Implementation Guidance
Organizations implementing our framework should consider several practical factors to maximize effectiveness. Initial deployment can begin immediately using existing tools and processes. Most vulnerability scanners already collect CVSS scores, requiring only the addition of KEV and EPSS data lookups. Simple scripts or spreadsheet formulas can implement the decision logic without requiring specialized security orchestration platforms.
Operational integration should align with existing patch management cycles. Critical Priority vulnerabilities (KEV-listed with CVSS $\geq$ 7.0) warrant emergency change procedures, while High Priority vulnerabilities can follow accelerated but planned maintenance windows. The Monitor category provides valuable input for threat hunting and detection engineering teams who can develop specific monitoring for exploited but lower-severity vulnerabilities.
Resource allocation becomes more strategic with clear prioritization tiers. Security teams can confidently defer approximately 95% of vulnerabilities to standard patching cycles, focusing specialized resources on the  5% with evidence of exploitation threat. This dramatic workload reduction enables deeper analysis of truly critical vulnerabilities, including understanding attack patterns, validating compensating controls, and ensuring complete remediation.
6.3 Evolution of Threat Intelligence Landscape
Since our data collection period, the vulnerability management landscape has continued to evolve rapidly. EPSS models undergo regular retraining with improved feature engineering and expanded threat intelligence feeds. KEV catalog growth has accelerated, with CISA adding vulnerabilities more frequently as threat intelligence sharing improves. Commercial vendors have introduced new prioritization frameworks attempting to address limitations we identify.